Circular No. 6541

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1636–53 W. Zhang, Goddard Space Flight Center (GSFC); I. Lapidus, Univer-

sity of Cambridge; J. H. Swank and N. E. White, GSFC; and L. Titarchuk, GSFC and George Mason University, report: “Four type-I x-ray bursts from 4U 1636–53 were detected during 1996 Dec. 25–1997 Jan. 10 with the RXTE. Each burst has nearly-coherent brightness oscillations at frequency 581 Hz, with fractional rms amplitudes ranging from 2.5 ± 0.3 to 8.7 ± 0.5 percent in the RXTE/PCA 2–60-keV band (stronger bursts having lower fractional rms amplitudes). No similar oscillations were detected at any other frequencies during these bursts. Also detected are two QPOs with frequencies in the ranges 840–920 and 1150–1200 Hz in the persistent flux; when both are present at the same time, their frequency difference appears to be a constant of 255 ± 25 Hz, similar to that reported on IAUC 6428. The 581 Hz differs from the frequency difference of the two simultaneous QPOs in the persistent flux. Considering its stability from burst to burst and its level of coherence, we think that 581 Hz, or possibly 290 Hz (although we have not observed any significant excess power for it), is most likely the neutron-star spin frequency. This is a second case, after Sco X-1 (IAUC 6424), where the frequency difference of the two QPOs in the persistent flux is not necessarily a valid measure of the neutron-star spin frequency, as a simple beat-frequency model would require. This is in contrast to the case of 4U 1728–34 (IAUC 6387), where the frequency difference and the frequency found during bursts are the same (namely, 363 Hz).”

COMET 29P/SCHWASSMANN-WACHMANN 1 Total CCD magnitude estimates: Jan. 11.77 UT, 13.2 (A. Nakamura,

Kuma, Ehime, Japan, 0.60-m reflector); 13.82, 13.4 (Nakamura); 16.38, 13.14 (I. Ferrin, National Observatory of Venezuela, 1-m Schmidt telescope; magnitude of inner 2′ of coma); 19.83, 13.0 (Nakamura).

COMET 81P/WILD 2 Total visual magnitude estimates: 1996 Dec. 22.08 UT, 11.5 (A. Baran-

sky, Kiev, Ukraine, 0.20-m reflector); 1997 Jan. 3.28, 11.1 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 5.26, 10.5 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 13.15, 10.7 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 18.12, 10.8 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1997 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green